# Linear Discriminant Regularized Regression

### v1

@@ A.2.1 Two key lemmas used in the proof of Theorem 2
Under model (1.1) and (1.3) with 1, assume $L\log(n\vee p)\leq n$. With probability at least $1-n^{-1}$, one has
$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbb{X}}_{j}^{\top}({\mathbb{Y}}_{\ell}-{\mathbb{X}}B_{\ell})\right|\lesssim\left(\sqrt{\|\Sigma_{w}\|_{\infty}}+\|{\mathbb{M}}\|_{\infty}\right)\sqrt{\log(n\vee p)\over nL}.$
@@
${1\over n}{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)$ $={1\over n}{\mathbb{X}}^{\top}{\mathbb{Y}}-\Sigma_{XY}-\widehat{\Sigma}B+\Sigma B$ (A.9)
We first analyze the term $(\widehat{\Sigma}-\Sigma)B$. Pick any $j\in[p]$, $\ell\in[L]$ and fix $t\geq 0$. We observe from Lemma 15 and the inequality $L\leq n$ that $\mathcal{E}_{\pi}$ holds with probability at least $1-2n^{-1}$. Next, we invoke Lemma 13 with $v_{1}={\bm{e}}_{j}$, $v_{2}=B_{\ell}$, $t=C\log(n\vee p)$ and we use (B.5), (B.6) and (B.7) in Lemma 9 to derive that the inequalities
@@
$\quad+\|{\mathbb{M}}_{j\cdot}\|_{\infty}\sqrt{B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{\mathbb{M}}_{j\cdot}\|_{\infty}\|{\mathbb{M}}^{\top}B_{\ell}\|_{\infty}{\log(n\vee p)\over n}$
$\quad+\sqrt{[\Sigma_{w}]_{jj}B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{\mathbb{M}}_{j\cdot}\|_{2}\sqrt{B_{\ell}^{\top}\Sigma_{w}B_{\ell}}\sqrt{\log(n\vee p)\over nL}$
$\lesssim\sqrt{[\Sigma_{w}]_{jj}+\|{\mathbb{M}}_{j\cdot}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}+\|{\mathbb{M}}_{j\cdot}\|_{\infty}{\log(n\vee p)\over n}$
$\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{\mathbb{M}}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$
@@
$\max_{j\in[p],\ \ell\in[L]}\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{\mathbb{M}}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (A.10)
@@
To bound the first term in (A.9), by recalling that $\Sigma_{XY}={\mathbb{M}}D_{\pi}$, we find
${1\over n}{\mathbb{X}}_{j}^{\top}{\mathbb{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}$ $={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{L}\left(X_{ij}Y_{i\ell}-{\mathbb{M}}_{j\ell}\pi_{\ell}\right)\mathbbm{1}\{Y_{i}={\bm{e}}_{\ell}\}$
$={1\over n}\sum_{i=1}^{L}\left(X_{ij}-{\mathbb{M}}_{j\ell}\right)\mathbbm{1}\{Y_{i}=\ell\}+\left({n_{\ell}\over n}-\pi_{\ell}\right){\mathbb{M}}_{j\ell}$
$={n_{\ell}\over n}(\widehat{\mathbb{M}}_{j\ell}-{\mathbb{M}}_{j\ell})+(\widehat{\pi}_{\ell}-\pi_{\ell}){\mathbb{M}}_{j\ell}.$
@@
$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbb{X}}_{j}^{\top}{\mathbb{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}\right|\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}\log(p\vee n)\over nL}+\|{\mathbb{M}}\|_{\infty}\sqrt{\log(n)\over nL}$ (A.11)
with probability $1-n^{-1}$. Combining (A.10) and (A.11) completes the proof. ∎
@@
Under model and (1.3) with Assumptions 1 & 2, assume $(s\vee L)\log(n\vee p)\leq n$ for some integer $1\leq s<p$. Then, on the event $\mathcal{E}_{\pi}$, with probability at least $1-(n\vee p)^{-2}$, the following holds uniformly over $v\in\Gamma_{s}$,
@@
$\sup_{v\in\Gamma_{s}^{\prime}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}}u^{\top}(\widehat{\Sigma}-\Sigma)u.$ (A.12)
@@
We will bound the right hand side (A.12) via the same arguments used to prove Lemma 13, except for the control of the term $\sum_{k}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$. Specifically, repeating the proof of Lemma 13 with $v_{1}=v_{2}=v$, for any $v\in\mathcal{N}$, gives that for any $t\geq 0$
$v^{\top}(\widehat{\Sigma}-\Sigma)v$ $\lesssim v^{\top}\Sigma_{w}v\left(\sqrt{t\over n}+{t\over n}\right)+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v+2{\|{\mathbb{M}}^{\top}v\|_{2}\over\sqrt{L}}\sqrt{v^{\top}\Sigma_{w}v}\sqrt{t\over n}.$
@@
${\|{\mathbb{M}}^{\top}v\|_{2}^{2}\over L}\lesssim v^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}v$
@@
$\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v\leq\max_{k\in[L]}{|\widehat{\pi}_{k}-\pi_{k}|\over\pi_{k}}v^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}v\lesssim v^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}v\sqrt{L\log(n)\over n}$
holds uniformly over $\mathcal{N}$. The proof is complete in view of (A.12). ∎
A.3 Proof of Theorem 3
@@
Define
$\delta\asymp\Delta_{\infty}\sqrt{\log(n)\over nL}+\sqrt{\Delta_{\infty}}\|{\mathbb{M}}\|_{\infty}\sqrt{s\log(n\vee p)\over nL}.$ (A.13)
We work on the event
$\mathcal{E}_{H}:=\left\{\|\widehat{H}-H\|_{\infty,1}\leq\delta\right\}$ (A.14)
which, according to Lemma 8, holds with probability at least $1-n^{-1}$. Moreover, on the event $\mathcal{E}_{H}$, we find that
$=\left[\lambda_{1}(\Omega)\right]^{-1}-\|\widehat{H}-H\|_{1,\infty}$ $\text{by writing }\Omega:=H^{-1}$
[…]
Hence, $\widehat{H}$ is invertible on the event $\mathcal{E}_{H}$. For simplicity, we write $\widehat{\Omega}=\widehat{H}^{-1}$. By definition,
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H\widehat{\Omega}-\mathbb{I}_{L})$ by $\text{by }H\Omega=\mathbb{I}_{L}$
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H-\widehat{H})\widehat{\Omega}$ by $\text{by }\widehat{H}\widehat{\Omega}=\mathbb{I}_{L}.$
Recall that (A.17) and pick any $\ell\in[L]$. We it then follows that
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\widehat{\Omega}_{\ell}\|_{2}+\|\Sigma_{w}^{1/2}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{2}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{2,\infty}\|\widehat{\Omega}_{\ell}\|_{1}+\max_{k\in[L]}\sqrt{B_{k}^{*\top}\Sigma_{w}B^{*}_{k}}~{}\|(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{1}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{2,\infty}\|\widehat{\Omega}_{\ell}\|_{1}+\sqrt{\Delta_{\infty}}\|\widehat{H}-H\|_{\infty,1}\|\widehat{\Omega}_{\ell}\|_{1}$ by $\text{by }B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}.$
To bound from above $\|\widehat{\Omega}_{\ell}\|_{1}$, we have that
$\|\widehat{\Omega}_{\ell}\|_{1}$ $\leq\|\Omega_{\ell}\|_{1}+\|(\widehat{\Omega}-\Omega){\bm{e}}_{\ell}\|_{1}$
$=\|\Omega_{\ell}\|_{1}+\|\Omega(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{1}$
$\leq\|\Omega_{\ell}\|_{1}+\|\Omega\|_{\infty,1}\|H-\widehat{H}\|_{\infty,1}\|\widehat{\Omega}_{\ell}\|_{1}.$
Since the event $\mathcal{E}_{H}$ and (3.3) ensure that
$\|\Omega\|_{\infty,1}\|H-\widehat{H}\|_{\infty,1}\leq\left(\min_{k\in[K]}{1\over\pi_{k}}+\|\Delta\|_{1,\infty}\right)\delta\lesssim\delta L\Delta_{\infty}\leq{1\over 2},$
we conclude, for any $\ell\in[L]$,
$\|\widehat{\Omega}_{\ell}\|_{1}\leq 2\|\Omega_{\ell}\|_{1}\lesssim L\Delta_{\infty}.$ (A.15)
It then follows that
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}$
$\lesssim\left(\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{2,\infty}+\delta\sqrt{\Delta_{\infty}}\right)L\Delta_{\infty}$
$\lesssim\Delta_{\infty}\|{\mathbb{M}}\|_{\infty}\sqrt{sL\log(n\vee p)\over n}+\sqrt{\Delta_{\infty}^{5}}\sqrt{L\log(n)\over n}+\Delta_{\infty}^{2}\|{\mathbb{M}}\|_{\infty}\sqrt{sL\log(n\vee p)\over n}$
$\lesssim\Delta_{\infty}^{5/2}\|{\mathbb{M}}\|_{\infty}\sqrt{sL\log(n\vee p)\over n},$ (A.16)
completing the proof. ∎
The following lemma provide upper bounds for $\|\widehat{H}-H\|_{\infty,1}$ under condition of Theorem 3.
@@
Under conditions of Theorem 3, with probability at least $1-n^{-1}$,
$\|\widehat{H}-H\|_{\infty,1}\lesssim\Delta_{\infty}\sqrt{\log(n)\over nL}+\sqrt{\Delta_{\infty}}\|{\mathbb{M}}\|_{\infty}\sqrt{s\log(n\vee p)\over nL}.$
@@
Recall that
@@
$\begin{split}\widehat{H}-H&=D_{\widehat{\pi}}-D_{\pi}-\left(\widehat{B}^{\top}\widehat{\Sigma}\widehat{B}-B^{\top}\Sigma B\right)\\ &=D_{\widehat{\pi}}-D_{\pi}-B^{\top}(\widehat{\Sigma}-\Sigma)B-(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\\ &\qquad-(\widehat{B}-B)^{\top}\widehat{\Sigma}B-B^{\top}\widehat{\Sigma}(\widehat{B}-B).\end{split}$ (A.17)
We first bound the $\|\cdot\|_{\infty,1}$ norm. By triangle inequality, we have
$\|\widehat{H}-H\|_{\infty,1}$ $\leq\|D_{\widehat{\pi}}-D_{\pi}\|_{\infty,1}+\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\infty,1}+\|(\widehat{B}-B)^{\top}\widehat{\Sigma}B\|_{\infty,1}$
$\quad+\|B^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\infty,1}+\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\infty,1}.$
We proceed to bound each term separately. On the event $\mathcal{E}_{\pi}$, we find that
$\|D_{\widehat{\pi}}-D_{\pi}\|_{\infty,1}\leq\max_{\ell\in[L]}|\widehat{\pi}_{\ell}-\pi_{\ell}|\lesssim\sqrt{\log(n)\over nL}.$ (A.18)
Next, pick any $k,\ell\in[L]$. We bound from above
$\textrm{I}=(\widehat{B}_{\ell}-B_{\ell})^{\top}\widehat{\Sigma}(\widehat{B}_{k}-B_{k})\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}^{2},$ (A.19)
$\textrm{II}=(\widehat{B}_{\ell}-B_{\ell})^{\top}\widehat{\Sigma}B_{k}\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}\|\widehat{\Sigma}^{1/2}B_{k}\|_{2},$
$\textrm{III}=B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}.$
Note that
$\|\widehat{\Sigma}^{1/2}B_{k}\|_{2}\leq\left(B_{k}^{\top}\Sigma B_{k}+B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}\right)^{1/2}.$
Meanwhile Lemma 14 ensures that, on the event $\mathcal{E}_{\pi}$ in (A.1),
$\textrm{III}\leq\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\infty}\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL}$ (A.20)
with probability at least $1-n^{-3}$. The above two bounds together with (B.5) in Lemma 9 give that
$\textrm{II}\lesssim{\sqrt{\Delta_{\infty}}\over L}\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}.$ (A.21)

### v2

@@ B.6.1 Two key lemmas used in the proof of Theorem 6
Under model (1.1) with 1, assume $L\log(n\vee p)\leq n$. With probability at least $1-n^{-1}$, one has
$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbf{X}}_{j}^{\top}({\mathbf{Y}}_{\ell}-{\mathbf{X}}B_{\ell})\right|\lesssim\left(\sqrt{\|\Sigma_{w}\|_{\infty}}+\|{M}\|_{\infty}\right)\sqrt{\log(n\vee p)\over nL}.$
@@
${1\over n}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)$ $={1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-\Sigma_{XY}-\widehat{\Sigma}B+\Sigma B$ (B.31)
We first analyze the term $(\widehat{\Sigma}-\Sigma)B$. Pick any $j\in[p]$, $\ell\in[L]$ and fix $t\geq 0$. We observe from Lemma 21 and the inequality $L\leq n$ that $\mathcal{E}_{\pi}$ holds with probability at least $1-2n^{-1}$. Next, we invoke Lemma 18 with $v_{1}={\bm{e}}_{j}$, $v_{2}=B_{\ell}$, $t=C\log(n\vee p)$ and we use (C.5), (C.6) and (C.7) in Lemma 14 to derive that the inequalities
@@
$\quad+\|{M}_{j\cdot}\|_{\infty}\sqrt{B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{M}_{j\cdot}\|_{\infty}\|{M}^{\top}B_{\ell}\|_{\infty}{\log(n\vee p)\over n}$
$\quad+\sqrt{[\Sigma_{w}]_{jj}B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{M}_{j\cdot}\|_{2}\sqrt{B_{\ell}^{\top}\Sigma_{w}B_{\ell}}\sqrt{\log(n\vee p)\over nL}$
$\lesssim\sqrt{[\Sigma_{w}]_{jj}+\|{M}_{j\cdot}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}+\|{M}_{j\cdot}\|_{\infty}{\log(n\vee p)\over n}$
$\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$
@@
$\max_{j\in[p],\ \ell\in[L]}\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (B.32)
@@
To bound the first term in (B.31), by recalling that $\Sigma_{XY}={M}D_{\pi}$, we find
${1\over n}{\mathbf{X}}_{j}^{\top}{\mathbf{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}$ $={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{L}\left(X_{ij}Y_{i\ell}-{M}_{j\ell}\pi_{\ell}\right)\mathbbm{1}\{Y_{i}={\bm{e}}_{\ell}\}$
$={1\over n}\sum_{i=1}^{L}\left(X_{ij}-{M}_{j\ell}\right)\mathbbm{1}\{Y_{i}=\ell\}+\left({n_{\ell}\over n}-\pi_{\ell}\right){M}_{j\ell}$
$={n_{\ell}\over n}(\widehat{M}_{j\ell}-{M}_{j\ell})+(\widehat{\pi}_{\ell}-\pi_{\ell}){M}_{j\ell}.$
@@
$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbf{X}}_{j}^{\top}{\mathbf{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}\right|\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}\log(p\vee n)\over nL}+\|{M}\|_{\infty}\sqrt{\log(n)\over nL}$ (B.33)
with probability $1-n^{-1}$. Combining (B.32) and (B.33) completes the proof. ∎
@@
Under model (1.1) with Assumptions 1 & 5, assume $(s\vee L)\log(n\vee p)\leq n$ for some integer $1\leq s<p$. Then, on the event $\mathcal{E}_{\pi}$, with probability at least $1-(n\vee p)^{-2}$, the following holds uniformly over $v\in\Gamma_{s}$,
@@
$\sup_{v\in\Gamma_{s}^{\prime}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}}u^{\top}(\widehat{\Sigma}-\Sigma)u.$ (B.34)
@@
We will bound the right hand side (B.34) via the same arguments used to prove Lemma 18, except for the control of the term $\sum_{k}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$. Specifically, repeating the proof of Lemma 18 with $v_{1}=v_{2}=v$, for any $v\in\mathcal{N}$, gives that for any $t\geq 0$
$v^{\top}(\widehat{\Sigma}-\Sigma)v$ $\lesssim v^{\top}\Sigma_{w}v\left(\sqrt{t\over n}+{t\over n}\right)+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v+2{\|{M}^{\top}v\|_{2}\over\sqrt{L}}\sqrt{v^{\top}\Sigma_{w}v}\sqrt{t\over n}.$
@@
${\|{M}^{\top}v\|_{2}^{2}\over L}\lesssim v^{\top}{M}D_{\pi}{M}^{\top}v$
@@
$\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v\leq\max_{k\in[L]}{|\widehat{\pi}_{k}-\pi_{k}|\over\pi_{k}}v^{\top}{M}D_{\pi}{M}^{\top}v\lesssim v^{\top}{M}D_{\pi}{M}^{\top}v\sqrt{L\log(n)\over n}$
holds uniformly over $\mathcal{N}$. The proof is complete in view of (B.34). ∎
B.7 Proof of Theorem 8: the reduced-rank estimator of $B$
@@
Let $C>0$ be some constant to be specified later. We work on the event
$\mathcal{E}_{\lambda}^{\prime}=\left\{{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}\leq\sqrt{\lambda/C}\right\}$
intersecting with
$\mathcal{E}_{X}=\left\{\lambda_{p}(\widehat{\Sigma})\geq{1\over 2}\lambda_{p}(\Sigma)\geq{c\over 2}\right\}.$
Here $c$ is defined in 3. Lemma 20, Lemma 13 and 3 together with (3.6) ensure that $\mathcal{E}_{X}\cap\mathcal{E}_{\lambda}^{\prime}$ hold with probability at least $1-n^{-1}$. Write $\widehat{r}={\rm rank}(\widehat{B})$ and $r={\rm rank}(B)$. By the optimality of $\widehat{B}$, we have
${1\over n}\|{\mathbf{Y}}-{\mathbf{X}}\widehat{B}\|_{F}^{2}+\lambda\widehat{r}$ $~{}\leq~{}{1\over n}\|{\mathbf{Y}}-{\mathbf{X}}B\|_{F}^{2}+\lambda r.$
Working out the squares gives
${1\over n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ $\leq{2\over n}|\langle{\mathbf{Y}}-{\mathbf{X}}B,{\mathbf{X}}(\widehat{B}-B)\rangle|+\lambda(r-\widehat{r})$
$\leq 2\|\widehat{B}-B\|_{*}{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}+\lambda(r-\widehat{r})$
$\leq 2\sqrt{r+\widehat{r}}~{}\|\widehat{B}-B\|_{F}\sqrt{\lambda/C}+\lambda(r-\widehat{r})$ by $\text{by $\mathcal{E}_{\lambda}^{\prime}$}.$
Using the properties on $\mathcal{E}_{X}$, we find
$\|\widehat{B}-B\|_{F}^{2}$ $\leq{2\over c}2\sqrt{\widehat{r}+r}\|\widehat{B}-B\|_{F}\sqrt{\lambda/C}+{2\over c}\lambda(r-\widehat{r})$
$\leq{8\lambda\over c^{2}C}(\widehat{r}+r)+{1\over 2}\|\widehat{B}-B\|_{F}^{2}+{2\over c}\lambda(r-\widehat{r})$ by $2xy\leq x^{2}/2+2y^{2}$
$={1\over 2}\|\widehat{B}-B\|_{F}^{2}+{4\over c}\lambda r$ for $\text{for $C=4/c$}.$
In the same way, on the event $\mathcal{E}_{X}$,
${1\over n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ $\leq\sqrt{{8\over cC}\lambda(r+\widehat{r})}~{}{1\over\sqrt{n}}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}+\lambda(r-\widehat{r})$
$\leq\frac{1}{2n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}+2\lambda r$ for $\text{for $C=4/c$}.$
This implies the rates for both $\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ and $\|\widehat{B}-B\|_{F}^{2}$. They imply in turn the rates for $\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}$ and $\|\widehat{B}-B\|_{2,1}$ invoking the inequalities
$\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}^{2}$ $\leq\|\Sigma^{1/2}(\widehat{B}-B)\|_{F}^{2}\leq{2\over c}\|\widehat{B}-B\|_{F}^{2}$ by (2.6)
and
$\|\widehat{B}-B\|_{1,2}^{2}$ $\leq{p}\|\widehat{B}-B\|_{F}^{2}.$
This completes the proof. ∎
@@
Under model (1.1) and 1, assume $(p+L)\log(n\vee p)\leq n$. With probability at least $1-2n^{-2}$, we have
${1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}\lesssim\sqrt{\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})}\sqrt{(p+L)\log(n)\over nL}.$
@@
By (B.31), it suffices to bound from above
@@
$\textrm{I}=\left\|(\widehat{\Sigma}-\Sigma)B\right\|_{\rm op},\qquad\textrm{II}=\left\|{1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-\Sigma_{XY}\right\|_{\rm op}=\left\|{1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-{M}D_{\pi}\right\|_{\rm op}.$
For I, by invoking Lemma 20, we find with probability at least $1-n^{-2}$,
I $\leq\|\Sigma^{1/2}\|_{\rm op}\|\Sigma^{-1/2}(\widehat{\Sigma}-\Sigma)\Sigma^{-1/2}\|_{\rm op}\|\Sigma^{1/2}B\|_{\rm op}$
$\lesssim\sqrt{\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})}\sqrt{(p+L)\log(n)\over nL}$
where in the last step we also use
$\|\Sigma\|_{\rm op}=\|\Sigma_{w}^{1/2}(\mathbf{I}_{p}+\Sigma_{w}^{-1/2}{M}D_{\pi}{M}^{\top}\Sigma_{w}^{-1/2})\Sigma_{w}^{1/2}\|_{\rm op}\leq\|\Sigma_{w}\|_{\rm op}\left(1+{\Delta_{\rm op}\over L}\right)\leq\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})$
as well as
$\|\Sigma^{1/2}B\|_{\rm op}^{2}=\|B^{\top}\Sigma B\|_{\rm op}=\|D_{\pi}{M}^{\top}B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L(L+\Delta_{\rm op})}\leq{1\over L}$
from Lemma 14.
Regarding II, by standard discretization argument, it suffices to bound from above
$\max_{v\in\mathcal{N}_{p}(1/3),u\in\mathcal{N}_{L}(1/3)}v^{\top}\left({1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-{M}D_{\pi}\right)u$
with $\mathcal{N}_{p}(1/3)$ and $\mathcal{N}_{L}(1/3)$ being the $(1/3)$-epsilon net of $\mathcal{S}_{p}$ and $\mathcal{S}_{m}$, respectively. Fix any $v\in\mathcal{N}_{p}(1/3)$ and $u\in\mathcal{N}_{L}(1/3)$. Observe that
$v^{\top}\left({1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-{M}D_{\pi}\right)u$ $=v^{\top}(\widehat{M}D_{\widehat{\pi}}-{M}D_{\pi})u$
$=v^{\top}(\widehat{M}-{M})D_{\widehat{\pi}}u+v^{\top}{M}\left(D_{\widehat{\pi}}-D_{\pi}\right)u$
$=\sum_{k=1}^{L}u_{k}{n_{k}\over n}v^{\top}(\widehat{\mu}_{k}-\mu_{k})+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})u_{k}v^{\top}\mu_{k}.$
Invoking Lemma 15 and Lemma 16 yields that, for any $t\geq 0$, with probability $1-4e^{-t/2}$,
II $\lesssim\|u\|_{2}\sqrt{tv^{\top}\Sigma_{w}v\over n}\sqrt{\max_{k}n_{k}\over n}+\sqrt{t\sum_{k=1}^{m}\pi_{k}u_{k}^{2}(v^{\top}\mu_{k})^{2}\over n}+{t\|u\|_{\infty}\|v^{\top}{M}\|_{\infty}\over n}$
$\lesssim\sqrt{t\|\Sigma_{w}\|_{\rm op}\over nL}+\sqrt{t\|\Sigma_{w}\|_{\rm op}\Delta_{\infty}\over nL}+\sqrt{\|\Sigma_{w}\|_{\rm op}\Delta_{\infty}}{t\over n}$
where in the last step we use
$\|v^{\top}{M}\|_{\infty}=\|v^{\top}\Sigma_{w}^{1/2}\|_{2}\|\Sigma_{w}^{-1/2}{M}\|_{2,\infty}\leq\|\Sigma_{w}\|_{\rm op}^{1/2}\sqrt{\Delta_{\infty}}.$